For the reaction: N2 (g) + C2H2 (g)  2 HCN (g), Kc is 2.3x10-4 at 300 ºC. What are the equilibrium concentrations of all species if the initial concentration of HCN was 5.0 M?
Answer #1

ICE Table:

[N2] [C2H2] [HCM] initial 5.0 change +1x +1x -2x equilibrium +1x +1x 5.0-2x

Equilibrium constant expression is

Kc = [HCN]^2/[N2]*[C2H2]

2.3*10^-4 = (5-2*x)^2/(1*x)^2

sqrt(2.3*10^-4) = (5-2*x)/(1*x)

0.0151658 = (5-2*x)/(1*x)

1.517*10^-2*x = 5-2*x

-5 + 2.015*x = 0

x = 2.48119

At equilibrium:

[N2] = +1x = +1*2.48119 = 2.48119 M

[C2H2] = +1x = +1*2.48119 = 2.48119 M

[HCN] = 5.0-2x = 5.0-2*2.48119 = 0.03763 M

Answer:

[N2] = 2.48 M

[C2H2] = 2.48 M

[HCN] = 0.0376 M